# VAEmo: Efficient Representation Learning for Visual-Audio Emotion with Knowledge Injection

### v1

@@ 4.2. Implementation Details
Model Configuration. Our VAEmo framework employs a ViT-Small architecture. The model maintains a consistent embedding dimension of 512 across all components with 8 attention heads. Our streamlined architecture consists of a 10-layer feature transformation encoder followed by a 2-layer fusion encoder , resulting in a total parameter count of 39M—significantly smaller than comparable models like HiCMAE-B (81M) and AV-HuBERT variants (103M).
Training Details. Our training proceeds in two distinct stages. For Stage 1 Pre-training, we pretrain on the full VoxCeleb2 dataset using 8 NVIDIA A800 GPUs with a batch size of 170 per GPU. We use the AdamW optimizer with a base learning rate of 1.2e-3 and a cosine learning rate schedule. This stage involves 100 epochs of training, requiring approximately 48 hours to complete. For Stage 2 Knowledge Injection, to incorporate emotional knowledge, we randomly sample 10% of the VoxCeleb2 dataset (approximately 100K samples) and generate emotion-focused captions using our MLLM pipeline. To prevent catastrophic forgetting, we employ LayerNorm tuning (Zhao et al., [n. d.]) where only LayerNorm parameters are updated while keeping other weights frozen. This stage involves 10 epochs of training on the captioned subset.
4.3. Comparison with State-of-the-Art Approaches
@@
On the lab-controlled categorical emotion recognition dataset CREMA-D, our method achieves 85.71% UAR and 85.68% WAR, setting a new benchmark that surpasses larger models like HiCMAE-B (Sun et al., 2024a) (84.91% UAR, 84.89% WAR). On MSP-IMPROV, we obtain 66.12% UAR and 76.79% WAR, outperforming the previous best approach AW-HuBERT (Tran et al., 2023) (65.72% UAR, 71.80% WAR) by +0.40% UAR and +4.99% WAR while using 62% fewer parameters. These results demonstrate our model’s exceptional capability in controlled recording environments.
Our method also excels at continuous emotion recognition on the Werewolf-XL dataset, achieving 39.80%, 69.72%, and 52.87% PCC for arousal, valence, and dominance dimensions, respectively. The improvements compared to HiCMAE (Sun et al., 2024a) are particularly substantial for arousal (+6.06%) and dominance (+12.21%), indicating our approach’s effectiveness at capturing subtle emotional intensity variations. These results collectively demonstrate that architectural simplicity coupled with our two-stage knowledge injection paradigm creates a powerful foundation for emotion recognition across diverse settings—from challenging in-the-wild scenarios to controlled laboratory environments, and across both categorical and dimensional emotion paradigms.
Furthermore, as shown in Fig. 4, our proposed model, VAEmo (39M parameters), outperforms other state-of-the-art approaches on the MAFW dataset with a significantly lower parameter count. Specifically, VAEmo achieves a WAR of 58.91%, while models such as HiCMAE-B (81M parameters) and S4D (101M parameters) perform at 56.17% and 58.44%, respectively. This demonstrates that VAEmo achieves better performance with a significantly lower parameter count compared to these approaches.
@@ 4.4. Ablation Studies
To better understand the impact of our design choices, we conduct a series of ablation studies on the MAFW, MSP-IMPROV, and Werewolf-XL datasets.
@@
Model Architecture Choices. To better understand our architectural design decisions, we compare single-stream versus dual-stream approaches. For fair comparison, all experiments were conducted using 100 epochs of Stage 1 training. As shown in Tab. 9, our single-stream architecture achieves comparable performance while using 45% fewer parameters (39M vs. 71M). The dual-stream model essentially represents a simplified version of HiCMAE without hierarchical connections and cross-attention interactions. The single-stream model’s efficiency comes from sharing parameters across modalities, creating a unified parameter space for both audio and visual processing. This shared space enables tighter integration where features are processed through identical transformations. Despite using only simple token concatenation for fusion, the model effectively captures cross-modal emotional correlations. This reveals that for emotion recognition, parameter sharing from early processing stages can lead to better generalization, even with simpler fusion approaches. While computational costs remain similar between approaches due to the single-stream model processing longer concatenated sequences, the reduced parameter count makes the single-stream architecture more deployment-friendly.
@@ 5. Conclusion
In this paper, we presented VAEmo, an efficient framework for audio-visual emotion representation learning that addresses the critical mismatch between general and emotion-specific representations. Through a carefully orchestrated two-stage pre-training approach on unlabeled audio-visual data, combined with our carefully designed emotional knowledge injection paradigm, our model effectively bridges the semantic gap between general audio-visual perception and emotion-specific understanding. Extensive experiments across multiple downstream benchmarks demonstrate the strong generalization capabilities of our learned representations, achieving state-of-the-art fine-tuning performance while maintaining a significantly more compact architecture than previous approaches.
@@ Acknowledgements.
The computation is completed on the HPC Platform of Hefei University of Technology.References

### v2

@@ 4.2. Implementation Details
Model Configuration. Our VAEmo framework employs a ViT-Small architecture. The model maintains a consistent embedding dimension of 512 across all components with 8 attention heads. Our streamlined architecture consists of a feature transformation encoder which is 10-layer transformer followed by a fusion encoder which is 2-layer transformer, resulting in a total parameter count of 39M—significantly smaller than comparable models like HiCMAE-B (81M) and AV-HuBERT variants (103M).
Training Details. Our training proceeds in two distinct stages. For Stage 1 Pre-training, we train the model on the full VoxCeleb2 dataset using 8 NVIDIA A800 GPUs with a batch size of 170 per GPU. We use the AdamW optimizer with a base learning rate of 1.2e-3 and a cosine learning rate schedule. This stage involves 100 epochs of training, requiring approximately 48 hours to complete. For Stage 2 Knowledge Injection, to incorporate emotional knowledge, we randomly sample 10% of the VoxCeleb2 dataset (approximately 100K samples) and generate emotion-focused captions using our MLLM pipeline. To prevent catastrophic forgetting, we employ LayerNorm tuning (Zhao et al., [n. d.]) where only LayerNorm parameters are updated while keeping other weights frozen. This stage involves 10 epochs of training on the captioned subset.
4.3. Comparison with State-of-the-Art Methods
@@
On the lab-controlled categorical emotion recognition dataset CREMA-D, our method achieves 85.71% UAR and 85.68% WAR, setting a new benchmark that surpasses larger models like HiCMAE-B (Sun et al., 2024a) (84.91% UAR, 84.89% WAR). On MSP-IMPROV, we obtain 66.12% UAR and 76.79% WAR, outperforming the previous best approach AW-HuBERT (Tran et al., 2023) (65.72% UAR, 71.80% WAR) by +0.40% UAR and +4.99% WAR while using 62% fewer parameters. These results demonstrate our model’s exceptional capability in controlled recording environments. Our method also excels at continuous emotion recognition on the Werewolf-XL dataset, achieving 39.80%, 69.72%, and 52.87% PCC for arousal, valence, and dominance dimensions, respectively. The improvements compared to HiCMAE (Sun et al., 2024a) are particularly substantial for arousal (+6.06%) and dominance (+12.21%), indicating our approach’s effectiveness at capturing subtle emotional intensity variations. These results collectively demonstrate that architectural simplicity coupled with our two-stage knowledge injection paradigm creates a powerful foundation for emotion recognition across diverse settings.
Furthermore, as shown in Fig. 4, our proposed model, VAEmo (39M parameters), outperforms other state-of-the-art approaches on the MAFW dataset with a significantly lower parameter count. Specifically, VAEmo achieves a WAR of 58.91%, while models such as HiCMAE-B (81M parameters) and S4D (101M parameters) perform at 56.17% and 58.44%, respectively.
@@ 4.4. Ablation Studies
Model Architecture Choices. To better clarify our architectural design decisions, we compare single-stream versus dual-stream approaches. For fair comparison, all experiments were conducted using 100 epochs of Stage 1 training. As shown in Tab. 9, our single-stream architecture achieves comparable performance while using 45% fewer parameters (39M vs. 71M). The dual-stream model essentially represents a simplified version of HiCMAE without hierarchical connections and cross-attention interactions. The single-stream model’s efficiency comes from sharing parameters across modalities, creating a unified parameter space for audio and visual processing. The shared space enables tighter integration where features are processed through identical transformations. This reveals that for emotion recognition, parameter sharing from early processing stages can lead to better generalization, even with simpler fusion methods. Although computation costs are similar, the smaller parameter size makes single-stream model easier to deploy.
@@ 5. Conclusion
In this paper, we presented VAEmo, an efficient framework for audio-visual emotion representation learning that addresses the critical mismatch between general and emotion-specific representations. Through a carefully orchestrated two-stage pre-training approach on unlabeled audio-visual data, combined with our carefully designed emotional knowledge injection paradigm, our model effectively bridges the semantic gap between general audio-visual perception and emotion-specific understanding. Extensive experiments across multiple downstream benchmarks demonstrate that the VAEmo achieves state-of-the-art performance with strong generalization and a compact architecture.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China (NSFC, No. 62202139, No. 62172138) and the Fundamental Research Funds for the Central Universities (No. JZ2025HGTB0226, No. JZ2024HGTG0310). The computation is completed on the HPC Platform of Hefei University of Technology.References
@@
- Hurst et al. (2024) Aaron Hurst, Adam Lerer, Adam P Goucher, Adam Perelman, Aditya Ramesh, Aidan Clark, AJ Ostrow, Akila Welihinda, Alan Hayes, Alec Radford, et al. 2024. Gpt-4o system card. arXiv preprint arXiv:2410.21276 (2024).